# LangForce: Bayesian Decomposition of Vision Language Action Models via Latent Action Queries

### v4

@@ 4.2 Experiments on RoboCasa
We evaluate our method on the RoboCasa GR1 Tabletop Manipulation Benchmark (Nasiriany et al., 2024), which consists of 24 diverse manipulation tasks. These tasks feature complex interactions with articulated objects and varied geometries, exemplified by specific tasks like “PnPBottleToCabinetClose” and “PnPCanToDrawerClose”, as well as scenarios involving appliances like microwaves and toasters. For training, we utilize the Humanoid Robot Tabletop Manipulation subset from the PhysicalAI-Robotics-GR00T-X-Embodiment-Sim (Bjorck et al., 2025b) dataset. Apart from the dataset, all experimental configurations are identical to those detailed in Section 4.1. To guarantee statistical significance, we evaluate each task using 50 independent trials and report the average success rate (Avg@50).
@@
The quantitative results on RoboCasa are presented in Table 2. Consistent with the empirical evidence presented in our motivation (Section 2), the VisionOnly baseline achieves a surprisingly high success rate of 44.7%, lagging only slightly behind the standard QwenGR00T baseline (47.8%). This observation reconfirms the prevalence of the vision shortcut in this benchmark, suggesting that a significant portion of tasks can be solved by relying solely on visual cues. However, LangForce breaks this performance ceiling, achieving a state-of-the-art average success rate of 50.4% and surpassing all competing baselines, including QwenOFT (48.8%) and Isaac-GR00T (47.6%). Crucially, our method demonstrates substantial gains in tasks where the vision-only policy falters. For instance, in the task “PnP Novel From Placemat To Plate”, LangForce achieves 70.0% success, significantly outperforming both the VisionOnly baseline (34.0%) and the standard QwenGR00T (48.0%). These results indicate that maximizing the LLR objective successfully forces the policy to extract and utilize task-specifying information from language, rather than settling for local optima based on visual shortcuts.
@@ 4.3 Preservation of General Capabilities
A prevalent view is that fine-tuning VLMs for robotic action generation (VLA training) diminishes the model’s foundational reasoning and multimodal understanding, potentially leading to a loss of general conversational abilities (Zhou et al., 2025; Xu et al., 2025; Hancock et al., 2025; Yu et al., 2026). ChatVLA (Zhou et al., 2025) attributes this to two factors: spurious forgetting, where robot training overwrites crucial visual-text alignments, and task interference, where competing control and understanding tasks degrade performance when trained jointly. We observe similar degradation in our baseline: as shown in Figure 4, the standard QwenGR00T model loses its ability to converse coherently even when prompted with pure text inputs. In contrast, LangForce remarkably preserves these capabilities when queried with language instructions.
It is crucial to note a nuanced distinction: while LangForce preserves normal text-only conversational ability, its general vision-language conversation (image + text inputs) can still degrade after VLA training. We hypothesize this is mainly because, in our training setup, the vision tower (and the multimodal fusion/projection layers) must be adapted for control, which can shift visual representations away from the pre-trained vision–language alignment manifold. Such specialization is expected for an embodied agent optimized for action. Importantly, LangForce helps prevent the text-only language behavior of the backbone from collapsing. In the baseline, the vision shortcut can render instruction tokens effectively redundant for control, weakening the training signal that encourages meaningful language processing and increasing drift in shared parameters, which manifests as failures even on pure text queries (Fig. 4 and Fig. 5). Conversely, our method’s LLR objective enforces a strong dependency on language. This acts as a regularizer that maintains the functional utility of instruction tokens, thereby preserving the backbone’s text-only conversational ability even as the visual modality is specialized for control. This preservation is of significant practical value: it ensures that the VLM backbone does not degenerate into a shallow feature mapper. By retaining its linguistic core, the agent preserves the potential for higher-level reasoning and generalization to novel instructions, which are the primary motivations for employing Foundation Models in robotics.
@@ 4.4 Ablation Studies
Effectiveness of Bayesian Decomposition. Comparing the full LangForce (63.5%) with the “+ Action Query” ablation (57.5%), we observe a significant performance boost (+6.0%). This indicates that while the architectural changes provide some benefit, the core improvement stems from our dual-branch Bayesian learning objective. By explicitly modeling and maximizing the pointwise mutual information (PMI) between instructions and actions, the model effectively overcomes the vision shortcut, validating the central hypothesis of this work.
Potential of Latent Action Queries. Even without the dual-branch definition, introducing Latent Action Queries (“QwenGR00T + Action Query”) improves upon the QwenGR00T baseline (55.2% $\to$ 57.5%). This suggests that Latent Action Queries function as a promising architectural inductive bias. Unlike standard approaches that feed full sequences of vision and language token embeddings into the action decoder, our query-based mechanism forces the VLM to compress and summarize task-relevant information into a compact set of latent tokens. From a computational perspective, this design is highly efficient. It decouples the complexity of the Diffusion Transformer (DiT) from the length of the VLM input context. Specifically, the complexity of condition processing in the DiT is reduced from $O(N^{2})$ (scaling with the massive number of vision-language tokens $N$) to $O(K^{2})$ (scaling with the small, constant number of query tokens $K$), thereby streamlining the action generation process.
@@ 5 Discussion
Rethinking Data Collection Strategies. Our experiments suggest that the deterministic mapping from visual scenes to language instructions ($H(\ell\mid v)\approx 0$) in goal-driven datasets is a significant factor contributing to the vision shortcut. To mitigate this, we hypothesize that a shift in data collection strategies could be beneficial. Prioritizing data collection in ambiguous scenarios—where the task cannot be inferred solely from the initial observation—might naturally increase the conditional entropy of language. By enriching datasets with scenes that support multiple valid tasks, models may be forced to rely more heavily on instructions for disambiguation.
Leveraging Human Data for Robustness. Recently, there has been growing interest in training robot models on large-scale human video data, such as HRDT (Bi et al., 2025), In-N-On (Cai et al., 2025), METIS (Fu et al., 2025), and PhysBrain (Lin et al., 2025). Unlike curated robot datasets, human activities are inherently multimodal and context-dependent; the same environment often hosts a wide variety of behaviors, potentially leading to a less sharp $p(\ell\mid v)$. We conjecture that injecting action knowledge from such rich human distributions might help mitigate the information collapse observed in robot-only datasets.
World Models as an Alternative Bayesian Formulation. Beyond the VLM framework focused on in this work, recent studies have also explored adapting World Models for VLA control, as seen in F1-VLA (Lv et al., 2025), Mantis (Yang et al., 2025), and InternVLA-A1 (contributors, 2026). From a theoretical perspective, these approaches can be viewed as an alternative instantiation of the Bayesian rule, specifically performing inverse dynamics on imagined futures. If we consider $v$ as a sequence of past frames $v_{\leq t}$, and treat the future state $v_{t+1}$ as a latent variable generated by the model (conditioned on $\ell$), the action inference can be expressed as:
@@
Here, the numerator $p(v_{t+1}\mid v_{\leq t},a,\ell)$ represents a world model (forward dynamics) predicting the future state. The term $p(a\mid v_{\leq t},\ell)$ serves as an action prior, and the denominator $p(v_{t+1}\mid v_{\leq t},\ell)$ represents the future prediction marginalized over actions. In this formulation, the policy execution involves first “imagining” a desired future $v_{t+1}$ consistent with $\ell$, and then inferring the optimal action $a$ via the equation above. Since world models are typically trained on vast amounts of video data, the predictive distribution (the numerator) is often rich and highly sensitive to the action $a$. We hypothesize that this sensitivity prevents the collapse of the numerator to the denominator. This suggests that world model-based architectures could offer another robust technical path toward solving the vision shortcut, which we plan to explore in future work.
@@ 6 Related Work
Vision-Language-Action Models. To bridge the gap between semantic understanding and physical control, Vision-Language-Action (VLA) models have emerged as a dominant paradigm. Early works like Octo (Team et al., 2024) established the transformer-based policy as a versatile initialization, utilizing diffusion heads to handle multimodal action distributions. Building on pre-trained VLMs, OpenVLA (Kim et al., 2024) and its optimized variant OpenVLA-OFT (Kim et al., 2025) fine-tune large language models for robotic control, demonstrating strong generalization. To further enhance action modeling, specialized architectures have been proposed. CogACT (Li et al., 2024b) and RDT-1B (Liu et al., 2025) employ dedicated diffusion transformers to capture complex action dynamics. Recent works explore advanced architectures and generation algorithms to better balance reasoning and control. The GR00T series (N1, N1.5, N1.6) (Bjorck et al., 2025b, a; GEAR-Team et al., 2025) utilizes a dual-system design, architecturally coupling a VLM for perception with a diffusion head for action generation. In parallel, the $\pi_{0}$ series ($\pi_{0}$, $\pi_{0.5}$, FAST) (Black et al., 2024, 2025; Pertsch et al., 2025) leverages Flow Matching as a superior generation objective, adopting a two-stage training paradigm that pre-trains the VLM backbone before fine-tuning. Specifically, FAST introduces discrete motion tokens to efficiently encode actions for the VLM, while $\pi_{0.5}$ incorporates subtask decomposition to enhance long-horizon planning. Other approaches like X-VLA (Zheng et al., 2025a) introduce embodiment-specific soft prompts to facilitate cross-embodiment generalization. By learning separate sets of embeddings for each data source, X-VLA effectively leverages heterogeneous robot data with minimal additional parameters. SpatialVLA (Qu et al., 2025) argues that spatial understanding is central to manipulation, introducing Ego3D Position Encoding and Adaptive Action Grids to inject 3D information and learn transferable spatial action knowledge. Then, VideoVLA (Shen et al., 2025) explores transforming video generation models into robot manipulators. By jointly predicting action sequences and future visual outcomes, it leverages the “visual imagination” of generative models to enhance generalization across novel tasks and objects. BayesVLA (Xu et al., 2025) also employs Bayesian decomposition to address language diversity limitations in VLA training. Using a CLIP-based encoder and diffusion transformer, it adopts a two-stage framework: the first stage trains a vision-conditioned prior on large-scale vision-action pairs to capture multimodal action distributions, while the second stage freezes the prior and injects language-aware tokens to refine policy conditioning on instructions. However, BayesVLA does not scale its training to large-scale VLMs, while other VLA approaches lack systematic solutions to the vision shortcut problem, where models ignore language instructions in goal-driven datasets.

### v5

@@ 4.3 Experiments on RoboCasa
We evaluate our method on the RoboCasa GR1 Tabletop Manipulation Benchmark (Nasiriany et al., 2024), which consists of 24 diverse manipulation tasks. These tasks feature complex interactions with articulated objects and varied geometries, exemplified by specific tasks like “PnPBottleToCabinetClose” and “PnPCanToDrawerClose”, as well as scenarios involving appliances like microwaves and toasters. For training, we utilize the Humanoid Robot Tabletop Manipulation subset from the PhysicalAI-Robotics-GR00T-X-Embodiment-Sim (Bjorck et al., 2025b) dataset. All other settings follow Section 4.1. To guarantee statistical significance, we evaluate each task using 50 independent trials and report the average success rate (Avg@50).
Table 4: Results of evaluating the VLA models with the GR1 robot in the RoboCasa Tabletop simulation environment. The results for Isaac-GR00T N1.5 and Isaac-GR00T N1.6 are sourced from the official Isaac-GR00T github repository (Bjorck et al., 2025b). The results for QwenGR00T, QwenPI, QwenOFT, QwenFAST are sourced from the official starVLA experiments (starVLA, 2025). We highlight the best results in bold and the second-best results with underline.
Task Isaac-GR00T N1.5 Isaac-GR00T N1.6 QwenGR00T +Qwen3VL QwenPI +Qwen3VL QwenOFT +Qwen3VL QwenFAST +Qwen3VL VisionOnly QwenGR00T LangForce +Qwen3VL PnP Bottle To Cabinet Close 54.0 51.5 46.0 26.0 30.0 38.0 70.0 72.0 PnP Can To Drawer Close 50.0 13.0 80.0 62.0 76.0 44.0 78.0 78.0 PnP Cup To Drawer Close 38.0 8.5 54.0 42.0 44.0 56.0 42.0 46.0 PnP Milk To Microwave Close 60.0 14.0 48.0 50.0 44.0 44.0 50.0 56.0 PnP Potato To Microwave Close 32.0 41.5 28.0 42.0 32.0 14.0 44.0 36.0 PnP Wine To Cabinet Close 38.0 16.5 46.0 32.0 36.0 14.0 40.0 46.0 PnP * to * Close (Avg) 45.3 24.2 50.3 42.3 43.7 35.0 54.0 55.7 PnP Novel From Cuttingboard To Basket 38.0 58.0 48.0 40.0 50.0 54.0 58.0 66.0 PnP Novel From Cuttingboard To Cardboardbox 46.0 46.5 40.0 46.0 40.0 42.0 26.0 40.0 PnP Novel From Cuttingboard To Pan 58.0 68.5 68.0 60.0 70.0 58.0 72.0 68.0 PnP Novel From Cuttingboard To Pot 62.0 65.0 52.0 40.0 54.0 58.0 50.0 48.0 PnP Novel From Cuttingboard To Tieredbasket 28.0 46.5 56.0 44.0 38.0 40.0 20.0 44.0 PnP Novel From Cuttingboard To * (Avg) 46.4 56.9 52.8 46.0 50.4 50.4 45.2 53.2 PnP Novel From Placemat To Basket 30.0 58.5 42.0 44.0 32.0 36.0 48.0 54.0 PnP Novel From Placemat To Bowl 60.0 57.5 44.0 52.0 58.0 38.0 32.0 62.0 PnP Novel From Placemat To Plate 56.0 63.0 48.0 50.0 52.0 42.0 34.0 52.0 PnP Novel From Placemat To Tieredshelf 36.0 28.5 18.0 28.0 24.0 18.0 16.0 24.0 PnP Novel From Placemat To * (Avg) 45.5 51.9 38.0 43.5 41.5 33.5 32.5 48.0 PnP Novel From Tray To Cardboardbox 52.0 51.5 38.0 34.0 44.0 28.0 50.0 50.0 PnP Novel From Tray To Plate 48.0 71.0 56.0 64.0 56.0 34.0 64.0 58.0 PnP Novel From Tray To Pot 60.0 64.5 50.0 44.0 62.0 46.0 52.0 62.0 PnP Novel From Tray To Tieredbasket 52.0 57.0 36.0 50.0 54.0 36.0 42.0 44.0 PnP Novel From Tray To Tieredshelf 32.0 31.5 16.0 28.0 30.0 16.0 16.0 22.0 PnP Novel From Tray To * (Avg) 48.8 55.1 39.2 44.0 49.2 32.0 44.8 47.2 PnP Novel From Plate To Bowl 58.0 57.0 60.0 52.0 60.0 52.0 26.0 54.0 PnP Novel From Plate To Cardboardbox 44.0 43.5 50.0 40.0 50.0 30.0 38.0 48.0 PnP Novel From Plate To Pan 60.0 51.0 54.0 36.0 66.0 48.0 44.0 54.0 PnP Novel From Plate To Plate 64.0 78.7 70.0 48.0 68.0 50.0 60.0 78.0 PnP Novel From Plate To * (Avg) 56.5 57.6 58.5 44.0 61.0 45.0 42.0 58.5 Average 48.2 47.6 47.8 43.9 48.8 39.0 44.7 52.6
@@
The quantitative results on RoboCasa are presented in Table 4. LangForce achieves a state-of-the-art average success rate of 52.6%, surpassing all competing baselines including QwenOFT (48.8%), Isaac-GR00T N1.5 (48.2%), and the direct baseline QwenGR00T (47.8%). Notably, our method substantially outperforms the VisionOnly baseline (44.7%), demonstrating the effectiveness of language grounding. Consistent with the empirical evidence presented in our motivation (Section 2.1), the VisionOnly baseline achieves a surprisingly high success rate of 44.7%, lagging only slightly behind the standard QwenGR00T baseline (47.8%). This observation reconfirms the prevalence of the vision shortcut in this benchmark, suggesting that a significant portion of tasks can be solved by relying solely on visual cues. However, LangForce breaks this performance ceiling, achieving a state-of-the-art average success rate of 52.6% and surpassing all competing baselines, including QwenOFT (48.8%), Isaac-GR00T N1.5 (48.2%), and Isaac-GR00T N1.6 (47.6%). Crucially, our method demonstrates substantial gains in tasks where the vision-only policy falters. For instance, in “PnP Novel From Placemat To Bowl”, LangForce reaches 62.0%, substantially surpassing both the VisionOnly baseline (32.0%) and QwenGR00T (44.0%). These results indicate that maximizing the LLR objective successfully forces the policy to extract and utilize task-specifying information from language, rather than settling for local optima based on visual shortcuts.
@@ 4.4 Preservation of General Capabilities
A prevalent concern in the field is that fine-tuning VLMs for robotic action generation (VLA training) often compromises the model’s foundational reasoning and multimodal understanding, leading to a degradation of general conversational abilities (Zhou et al., 2025; Xu et al., 2025; Hancock et al., 2025; Yu et al., 2026). ChatVLA (Zhou et al., 2025) attributes this to spurious forgetting, where robot training overwrites critical visual-text alignments, and task interference between control and understanding objectives.
We observe similar degradation in our baseline: as illustrated in Figure 4 and Figure 5, the standard QwenGR00T model loses the ability to converse coherently, even when prompted with pure text inputs. In stark contrast, LangForce remarkably preserves these capabilities when queried with language instructions. This is because our LLR objective enforces a strong dependency on language for action generation, preventing the linguistic representations from collapsing or becoming redundant.
It is important to qualify these findings with a nuanced distinction: while LangForce effectively preserves the model’s text-only conversational abilities, its general vision-language capabilities (involving both image and text inputs) may still experience some degradation following VLA training. We attribute this to our training setup, which requires the vision tower and multimodal projection layers to adapt for specific control tasks, inevitably shifting visual representations away from the original pre-trained alignment manifold. Such specialization appears to be a natural trade-off when optimizing an embodied agent for precise action execution.
Crucially, however, LangForce successfully prevents the catastrophic collapse of the backbone’s text-only language behavior. In standard baselines, the prevalence of vision shortcuts often renders instruction tokens effectively redundant for control. This redundancy weakens the gradient signal for language processing, leading to a drift in shared parameters that manifests as failures even on pure text queries (as seen in Fig. 4 and Fig. 5). In contrast, our LLR objective acts as a powerful regularizer by enforcing a strong, explicit dependency on language. This mechanism maintains the functional utility of instruction tokens, thereby preserving the backbone’s text-only conversational ability even as the visual modality specializes for control. This retention of the linguistic core is of significant practical value: it ensures that the VLM does not degenerate into a shallow feature mapper, but rather preserves its potential for higher-level reasoning and generalization—key motivations for employing Foundation Models in robotics.
@@ 4.5 Ablation Studies
Effectiveness of Bayesian Decomposition.
Comparing the full LangForce (63.5%) with the “+ Action Query” ablation (57.5%), we observe a significant performance boost (+6.0%). This indicates that while the architectural changes provide some benefit, the core improvement stems from our dual-branch Bayesian learning objective. By explicitly modeling and maximizing the pointwise mutual information (PMI) between instructions and actions, the model effectively overcomes the vision shortcut, validating the central hypothesis of this work. Additionally, we provide ablation studies on the hyperparameters $\lambda$ and $\beta$ in Appendix C.1.
@@
Potential of Latent Action Queries.
Even without the dual-branch definition, introducing Latent Action Queries (“+ Action Query”) improves upon the QwenGR00T baseline (55.2% $\to$ 57.5%). This suggests that Latent Action Queries function as a promising architectural inductive bias. Unlike standard approaches that feed full sequences of vision and language token embeddings into the action decoder, our query-based mechanism forces the VLM to compress and summarize task-relevant information into a compact set of latent tokens. From a computational perspective, this design is highly efficient. It decouples the complexity of the Diffusion Transformer (DiT) from the length of the VLM input context. Specifically, the complexity of condition processing in the DiT is reduced from $O(N^{2})$ (scaling with the massive number of vision-language tokens $N$) to $O(K^{2})$ (scaling with the small, constant number of query tokens $K$), thereby streamlining the action generation process. We also investigate the impact of the number of latent action queries on model performance. Detailed analysis and results are provided in Appendix C.2.
@@ 5 Discussion
Rethinking Data Collection Strategies.
Our experiments suggest that the deterministic mapping from visual scenes to language instructions ($H(\ell\mid v)\approx 0$) in goal-driven datasets is a significant factor contributing to the vision shortcut. To mitigate this, we hypothesize that a shift in data collection strategies could be beneficial. Prioritizing data collection in ambiguous scenarios—where the task cannot be inferred solely from the initial observation—might naturally increase the conditional entropy of language. By enriching datasets with scenes that support multiple valid tasks, models may be forced to rely more heavily on instructions for disambiguation. Consequently, we advocate for:
Prioritizing ambiguous scenarios during data collection to naturally compel models to rely on language for disambiguation.
Leveraging Human Data for Robustness.
Recently, there has been growing interest in training robot models on large-scale human video data, such as HRDT (Bi et al., 2025), In-N-On (Cai et al., 2025), METIS (Fu et al., 2025), and PhysBrain (Lin et al., 2025). Unlike curated robot datasets, human activities are inherently multimodal and context-dependent; the same environment often hosts a wide variety of behaviors, potentially leading to a less sharp $p(\ell\mid v)$. This observation highlights a critical opportunity for scaling up robot learning:
Leveraging human video data, with its rich behavioral diversity, to mitigate the information collapse typical of curated robot datasets.
World Models as an Alternative Bayesian Formulation.
Beyond the VLM framework focused on in this work, recent studies have also explored adapting World Models for VLA control, as seen in F1-VLA (Lv et al., 2025), Mantis (Yang et al., 2025b), and InternVLA-A1 (contributors, 2026). From a theoretical perspective, these approaches can be viewed as an alternative instantiation of the Bayesian rule, specifically performing inverse dynamics on imagined futures. If we consider $v$ as a sequence of past frames $v_{\leq t}$, and treat the future state $v_{t+1}$ as a latent variable generated by the model (conditioned on $\ell$), the action inference can be expressed as:
@@
Here, the numerator $p(v_{t+1}\mid v_{\leq t},a,\ell)$ represents a world model (forward dynamics) predicting the future state. The term $p(a\mid v_{\leq t},\ell)$ serves as an action prior, and the denominator $p(v_{t+1}\mid v_{\leq t},\ell)$ represents the future prediction marginalized over actions. In this formulation, the policy execution involves first “imagining” a desired future $v_{t+1}$ consistent with $\ell$, and then inferring the optimal action $a$ via the equation above. Since world models are typically trained on vast amounts of video data, the predictive distribution (the numerator) is often rich and highly sensitive to the action $a$. We hypothesize that this sensitivity prevents the collapse of the numerator to the denominator, offering a robust technical path for future research:
Exploring World Models as an alternative Bayesian formulation, where the sensitivity of forward dynamics inherently resists the vision shortcut.
@@ 6 Related Work
Vision-Language-Action Models. To bridge the gap between semantic understanding and physical control, Vision-Language-Action (VLA) models have emerged as a dominant paradigm. Early works like Octo (Team et al., 2024) established the transformer-based policy as a versatile initialization, utilizing diffusion heads to handle multimodal action distributions. Building on pre-trained VLMs, OpenVLA (Kim et al., 2024) and its optimized variant OpenVLA-OFT (Kim et al., 2025) fine-tune large language models for robotic control, demonstrating strong generalization.
To further enhance action modeling, specialized architectures have been proposed. CogACT (Li et al., 2024b) and RDT-1B (Liu et al., 2025) employ dedicated diffusion transformers to capture complex action dynamics. Recent works explore advanced architectures and generation algorithms to better balance reasoning and control. The GR00T series (N1, N1.5, N1.6) (Bjorck et al., 2025b, a; GEAR-Team et al., 2025) utilizes a dual-system design, architecturally coupling a VLM for perception with a diffusion head for action generation. In parallel, the $\pi_{0}$ series ($\pi_{0}$, $\pi_{0.5}$, FAST) (Black et al., 2024, 2025; Pertsch et al., 2025) leverages Flow Matching as a superior generation objective, adopting a two-stage training paradigm that pre-trains the VLM backbone before fine-tuning. Specifically, FAST introduces discrete motion tokens to efficiently encode actions for the VLM, while $\pi_{0.5}$ incorporates subtask decomposition to enhance long-horizon planning.
@@
Other approaches like X-VLA (Zheng et al., 2025a) introduce embodiment-specific soft prompts to facilitate cross-embodiment generalization. By learning separate sets of embeddings for each data source, X-VLA effectively leverages heterogeneous robot data with minimal additional parameters. SpatialVLA (Qu et al., 2025) argues that spatial understanding is central to manipulation, introducing Ego3D Position Encoding and Adaptive Action Grids to inject 3D information and learn transferable spatial action knowledge. Then, VideoVLA (Shen et al., 2025) explores transforming video generation models into robot manipulators. By jointly predicting action sequences and future visual outcomes, it leverages the “visual imagination” of generative models to enhance generalization across novel tasks and objects. TwinBrainVLA (Yu et al., 2026) mitigates catastrophic forgetting by coordinating a frozen generalist “Left Brain” for semantic understanding and a trainable specialist “Right Brain” for sensorimotor learning, effectively balancing high-level reasoning with low-level control. BayesVLA (Xu et al., 2025) also employs Bayesian decomposition to improve instruction following. It utilizes a two-stage framework: first training a vision-conditioned prior on large-scale vision-action pairs, and then freezing this prior to train a language-conditioned posterior. However, this design strictly requires a separated training process where the vision component must be frozen in the second stage. In contrast, LangForce enables single-stage, end-to-end training. Furthermore, BayesVLA has not open-sourced its code, which regrettably prevents a direct empirical comparison. Overall, other VLA approaches lack systematic solutions to the vision shortcut problem, where models ignore language instructions in goal-driven datasets.